A Kitten Is Given An Injection Of A 0.5% Solution. The Dose Rate Is 1mg/Kg/Day And The Kitten Weighs 0.5lg. What\'S The Volume Injected?

1
Expert's answer
2021-11-09T11:48:40-0500

"1\frac{mg}{kg}*0.5kg=0.5mg."

"x*0.5\%=0.5mg."

"x=\frac{0.5}{0.005}=100 ml."

The volume injected is 100 ml.